php?id 4954882 TBOOK 17 of 24 How many grams of sodium iodide, Nal, must be used to produce 67.4 g of iodine, 12? g Nal t ther
Answer #1

From the stoichiometry of the reaction

1 mol I2 produced from = 2 moles of NaI

253.81 g I2 produced from = 2*149.89 = 299.78 g of NaI

67.4 g I2 produced from = 299.78*67.4/253.81

= 79.61 g of NaI